# smile: Single-turn to Multi-turn Inclusive Language Expansion via ChatGPT for Mental Health Support

### v2

@@ 4.2 Analysis
[…]
Figure 3 presents the distribution of dialogue transformation among three methods. According to our analysis, we conclude that single-turn dialogues can be successfully rewritten into multi-turn dialogues.
@@
5 Dialogue Diversity
[…]
5.1 Semantic Features
[…]
To calculate the pairwise cosine similarity between every two distinct dialogues, this produces $\binom{500}{2}$ pairs of dialogues and their corresponding cosine values in each prompt method, including the reference point of sampled data from PsyQA.
We present the results in Figure 4, which demonstrates that the median of the smile method is significantly lower than those of the baseline methods. The smile method exhibits the most extensive semantic diversity, aligning closely with the sampled dialogue of PsyQA. However, it’s worth noting that the median of the smile method is larger than that of PsyQA*. The reason behind this may be the introduction of token distribution from ChatGPT.
Category Total Help-seeker Supporter
# Dialogues 55165 - -
# Utterances 1833856 693756 1140100
# Avg. turns per dialogue 10.4 - -
Avg. utterances per dialogue 33.2 12.6 20.7
Avg. length per utterance 27.9 26.1 28.9
Table 1: Statistics of the dialogue dataset, SmileChat.
6 Expert Evaluation
We elaborate on data quality analysis from two main professional perspectives: (1) the evaluation of dialogue quality by three professional counselors, and (2) alignment with the supporters’ strategy flow patterns in real-life counseling conversations (For brevity, we include this section in Appendix G.).
To further validate that the dialogue quality generated using the SMILE method is superior to that produced by the two baseline methods, we propose to recruit three professional psychological counselors and randomly sample 20 dialogues for each method. Each set comprises three dialogues generated from different methods. Three counselors assign scores (3, 2, and 1, where 3 is the best, and 1 is the worst) based on three metrics: assessing (Q1) the overall quality of the conversation, (Q2) the authenticity of the user’s stated issues, and (Q3) the professionalism of the counselor’s responses.
We present the results of an expert evaluation of dialogue quality in 20 sets of dialogues in Table 2. Each value in the table represents the average score from three professional counselors, and results show that the dialogue quality generated with the SMILE method is of higher quality than that generated by the two baseline methods. The intraclass correlation coefficients (James et al., 1984; Koo and Li, 2016), including nine coefficients for the three metrics, are all greater than 0.8, indicating almost perfect agreement.
Method Q1 Q2 Q3
standard 1.95 2.12 1.21
stand.${}_{\mathrm{DT}}$ 1.25 1.2 1.82
smile 2.8 2.68 2.97
Table 2: Evaluation of dialogue quality by three professional counselors.
7 SmileChat Dataset
Through the analysis of language transformation, dialogue diversity and expert evaluation, we conclude that the proposed method can generate a diverse and high-quality chat dataset. Therefore, we utilize the smile method to guide ChatGPT in generating all multi-turn conversations based on PsyQA one round, leading to a large-scale dialogue dataset.
7.1 Data Statistics
To ensure data quality, we impose stricter requirements on dialogue turns, retaining only dialogues with at least 5 turns. Thus, we compile a collection of 55,165 conversations, SmileChat. Table 1 presents the statistics of the collected corpus.
7.2 Dialogue Exemplars
Multi-turn dialogue examples generated with the standard and stand.${}_{\mathrm{DT}}$ methods are illustrated in Figure 13 and 15. Further, a dialogue generated by the smile method is shown in Figure 17.
8 Dialogue System
We aim to build a high-quality multi-turn chat dataset for mental health. Therefore, we also analyze the dialogue quality based on the performance of the dialogue system trained with SmileChat.
8.1 Task Formulation
Our collected dataset can be represented as $D=\{d_{1},d_{2},...,d_{n}\}$, where each $d_{j}$ represents a single multi-turn dialogue.
To train a dialogue system for mental health, we need to split each dialogue into several training sessions. Specifically, a sampled $t$-turn dialogue session can be represented as follows:
$d_{t}=\left\{u^{H}_{0},u^{S}_{1},u^{H}_{2},...,u^{H}_{t-1},u^{S}_{t}\right\}\sim D$ (3)
We build a dialogue model that can predict the supporter’s utterance $u^{S}_{t}$ based on the dialogue history $h_{t}=\left\{u^{H}_{0},u^{S}_{1},u^{H}_{2},...,u^{H}_{t-1}\right\}$. Our objective is to maximize the likelihood probability as follows:
[…]
where $L$ is the sequence length of $u^{S}_{t}$.
8.2 Experimental Setup
Baseline Model
To validate the dialogue quality of our collected dataset, we conduct a fine-tuning experiment on ChatGLM2-6B (Zeng et al., 2023). Moreover, we also add two prompt engineering baselines (Prompt A and B, for more details, see Appendix I).
Parameter-efficient Fine-tuning
To preserve the original capabilities of the model while adapting to downstream dialogue tasks and reducing computational costs, we employ Low-Rank Adaptation (LoRA, (Hu et al., 2021)) on all linear layers in the ChatGLM2-6B model for efficient fine-tuning.
Hyperparameters
We present the hyperparameters for constructing a dialogue model for mental health in Table 10. For additional details on the training and generation processes, please refer to Appendix H.
Category Total Help-seeker Supporter
# Dialogues 82 - -
# Utterances 7672 3627 4045
# Avg. turns per dialogue 74.8 - -
Avg. utterances per dialogue 93.6 44.2 49.3
Avg. length per utterance 24.6 29.7 20.1
Table 3: Statistics of test set, PsyTest.
Methods PPL ($\Downarrow$) METEOR ($\Uparrow$) BLEU-1 ($\Uparrow$) BLEU-2 ($\Uparrow$) BLEU-3 ($\Uparrow$) Rouge-L ($\Uparrow$) D-1 ($\Uparrow$) D-2 ($\Uparrow$) D-3 ($\Uparrow$)
Baseline 3.38 ${}^{\dagger}$${}^{\ddagger}$ ${}^{\dagger}$${}^{\ddagger}$ ${}^{\dagger}$${}^{\ddagger}$ ${}^{\dagger}$${}^{\ddagger}$ ${}^{\dagger}$${}^{\ddagger}$ ${}^{\dagger}$${}^{\ddagger}$ ${}^{\dagger}$${}^{\ddagger}$ ${}^{\dagger}$${}^{\ddagger}$
Baseline + Prompt A 3.43 ${}^{\dagger}$${}^{\ddagger}$ ${}^{\dagger}$${}^{\ddagger}$ ${}^{\dagger}$${}^{\ddagger}$ ${}^{\dagger}$${}^{\ddagger}$ ${}^{\dagger}$${}^{\ddagger}$ ${}^{\dagger}$${}^{\ddagger}$ ${}^{\dagger}$${}^{\ddagger}$ ${}^{\dagger}$${}^{\ddagger}$
Baseline + Prompt B 3.99 ${}^{\dagger}$${}^{\ddagger}$ ${}^{\dagger}$${}^{\ddagger}$ ${}^{\dagger}$${}^{\ddagger}$ ${}^{\dagger}$${}^{\ddagger}$ ${}^{\dagger}$${}^{\ddagger}$ ${}^{\dagger}$${}^{\ddagger}$ ${}^{\dagger}$${}^{\ddagger}$ ${}^{\dagger}$${}^{\ddagger}$
Fine-tuned 1.50 ${}^{\dagger}$${}^{\ddagger}$ ${}^{\dagger}$${}^{\ddagger}$ ${}^{\dagger}$${}^{\ddagger}$ ${}^{\dagger}$${}^{\ddagger}$ ${}^{\dagger}$${}^{\ddagger}$ ${}^{\dagger}$${}^{\ddagger}$ ${}^{\dagger}$${}^{\ddagger}$ ${}^{\dagger}$${}^{\ddagger}$
Table 4: Results of automatic evaluation in our compiled real-life test set. $\dagger$ denotes the tokenizer we use is THUDM/chatglm2-6b, while $\ddagger$ denotes the tokenizer we use is hfl/chinese-roberta-wwm-ext-large.
8.3 Evaluation
To better understand and assess the dialogue quality of SmileChat dataset, we propose to utilize real-life multi-turn counseling conversations. We develop an online mental health support platform that enables professional counselors to offer each client a free text-based counseling service, lasting approximately 50 minutes each time. This duration is widely recognized as a standard time setting in psychological counseling. Therefore, we compile a collection of 82 high-quality real-life counseling dialogues with data desensitization for model evaluation. We name this test set PsyTest, which contains 4045 utterances spoken by supporters, as shown in Table 3.
8.3.1 Automatic Evaluation
Metrics
To conduct automatic evaluation, the evaluation metrics we use consist of Perplexity (PPL) (Jelinek et al., 1977), METEOR (Banerjee and Lavie, 2005), BLEU-1/2/3 (Papineni et al., 2002), Rouge-L (Lin, 2004), and Distinct-1/2/3 (D-1/2/3) (Li et al., 2016).
Results
The results of the automatic evaluation, including 9 metrics, are presented in Table 4. Notably, the evaluated dialogues are based on real-world counseling data rather than generated dialogues, which excludes the influence stemming from ChatGPT. All automatic evaluation metrics we use indicate improved performance. Our results show that the model trained with SmileChat is effective and practical. Consequently, the automatic evaluation demonstrates that our collected dataset is of high quality.
8.3.2 Human Evaluation
Metrics
We conduct a pairwise human evaluation to study the model performance trained with our proposed dialogue corpus. Initially, we randomly sample 100 multi-turn dialogue sessions, each comprising a minimum of 9 turns selected from PsyTest. Each dialogue session concludes with the utterance spoken by the counselor (also referred to as the supporter). Subsequently, we obtain 100 generated responses from the MeChat model. Three professional counselors are then presented with a dialogue history and three randomly shuffled responses (baseline, fine-tuned, ground truth). They are tasked with selecting the optimal response for the dialogue history, considering aspects such as naturalness, human-like qualities, correctness, and appropriateness. The evaluation is conducted based on the ethical principles of psychologists and the code of conduct (Association et al., 2016).
Figure 5: Pairwise human evaluation results, comparing the fine-tuned model to the baseline and ground truth, are reported. We present the win and lose rates of each compared pair in 100 randomly sampled real-life dialogue sessions. Fleiss’ kappa (Fleiss et al., 1981) is used to measure the inter-rater agreement, and all values fall within moderate agreement with $0.5\leq\kappa\leq 0.6$.
Results
We employ majority voting to reach final decisions among three professional counselors. As depicted in Figure 5, the model trained with SmileChat demonstrates a significant performance improvement compared to the baseline model. Moreover, the responses generated by MeChat surpass the golden response to some extent in 36% of all dialogue sessions.
Case Study
We present a case study, as shown in Figure 6. First, during counseling conversations, the text should not be too long for the help-seeker to understand. Second, at the beginning of a conversation, a supporter generally will not challenge the help-seeker; instead, they will provide support. Third, response 1 is more human-like and natural. Therefore, response 1 is better than response 2 overall.
Figure 6: Case study. Counseling strategies used in the two responses are presented in parentheses. Strategies in green are supporting, while those in red are challenging.
9 Conclusion
This paper introduces smile, a simple yet effective solution for addressing the scarcity of multi-turn conversations in mental health. Through language transformation and diversity analysis, we confirm the feasibility and effectiveness of our approach, demonstrating that our method is transferable. We also conduct expert evaluations, demonstrating that the smile method exhibits high superiority. Thus, the proposed method enables the automatic creation of a large-scale, diverse, and high-quality dialogue corpus, SmileChat, consisting of 55,165 dialogues with an average of 10.4 turns. Both automatic and human evaluations using the PsyTest dataset, consisting of 82 real-life counseling dialogues, demonstrate that SmileChat significantly improves dialogue system performance in mental health. With the release of multi-turn dialogues (SmileChat), our dialogue model (MeChat), and an authentic test set (PsyTest), we contribute valuable resources to the research community.
Limitations
We release a large-scale, diverse, and high-quality multi-turn conversational dataset for mental health support, generated by rewriting single-turn conversations into multi-turn conversations using ChatGPT. Consequently, the dataset unavoidably incorporates ChatGPT’s model knowledge. Furthermore, we discuss how ChatGPT does not fully utilize the rich vocabulary and content of single-turn conversations, as reflected in the distinct-$n$ metric.
Additionally, there is inevitably a gap in counseling practicality between models trained on this synthetic dataset and those trained on 50-minute counseling conversations, which will be our future work. While dialogues generated by the three methods still exhibit significant gaps when compared to the Xingling (Oracle) dataset with respect to the flow patterns of supporters’ strategies, there exists a large space for us to improve in our future work.
Furthermore, there is currently no comprehensive dataset available for evaluating the conversational safety of mental health support models. Therefore, we have identified this as a limitation of our paper and intend to address it in future work.
Ethical Considerations
We present a detailed illustration for the section on ethical considerations, which includes Data Usage Agreement, Usage Policy, and Suggestions for Real-life Deployment. Furthermore, the section on suggestions for real-life deployment highlights four main aspects: Anonymity and Consent, Handling Sensitive Topics, Ensuring Client Confidentiality, and Preventing Misinformation.
Data Usage Agreement
PsyQA is a single-turn dialogue dataset collected from an online mental health support platform. Specifically, help-seekers submit a post containing their mental health states and issues, and many professional counselors write down their responses to user questions to assist help-seekers. Therefore, PsyQA is a high-quality Chinese dialogue related to mental health support in the form of one question mapping to multiple answers.
Following the data copyright guidelines formulated by PsyQA (Sun et al., 2021), we release the multi-turn dialogue corpus publicly available for research purposes only. If researchers wish to reproduce the multi-turn dialogues using PsyQA, they must sign an agreement with the original data owner. Accordingly, we release our datasets and models for research purposes, thus facilitating further advancement in the academic community.
Usage Policy
Our chatbot is trained with the SmileChat dataset end-to-end. As illustrated in our paper, SmileChat still exhibits a significant gap in the flow patterns of counseling strategies compared to real-life counseling dialogues. Access to SmileChat and Mechat should be strictly limited to research purposes and should NOT be utilized for real-world deployment, commercial purposes, or any other use beyond academic research. Researchers using SmileChat and MeChat should be aware of the dataset and model limitations and make efforts to acknowledge and/or mitigate them to the best of their ability.
Suggestions for Real-life Deployment
When deploying our chatbot on the server for research purposes, we recommend implementing safeguards to protect help-seekers and, to a large extent, mitigate potential physical risks.
Anonymity and Consent
We strongly recommend that researchers allow users to interact with the chatbot anonymously, especially when discussing sensitive issues. Additionally, researchers should clearly communicate the level of anonymity and obtain informed consent from users before collecting any personal information.
Handling Sensitive Topics
In real-life human-machine interactions, we strongly recommend constructing a classifier to identify user intent. When users discuss sensitive topics, if detected by the intent classifier, a "human-in-the-loop" system allows a human moderator or healthcare professional to intervene in complex or sensitive situations.
Furthermore, we recommend ensuring clear escalation paths for users who require immediate human assistance, such as in cases of suicide and self-injury.
Ensuring Client Confidentiality
Researchers should clearly disclose to users that they are interacting with a chatbot, not a human, setting appropriate expectations for the system’s capabilities and limitations.
Researchers should provide clear information on how data is handled, stored, and for what purposes.
Misinformation Prevention
Researchers should implement fact-checking mechanisms to verify the accuracy of information provided by the chatbot.
Researchers should clearly communicate the limitations of the chatbot’s knowledge and advise users to consult a human professional for critical decisions.
Appendix A Details of Data Cleaning
A.1 Automatic Cleaning
We employ a sequential data cleaning pipeline to swiftly replace words that are unsuitable to the conversation scenario. For example, both "楼主你" (literally thread starter you) and "楼主" (literally thread starter) should be replaced with "你" (you). However, it is necessary to perform the former replacement to avoid the repetition of "你" and the resulting "你你" (you-you).
A.1.1 Word List for Data Cleaning
To avoid the repetition of "你" (you) and the resulting "你你" (you-you) , we suggest to conduct a sequential word replacing pipeline. Figure 7 shows the word list for data cleaning and corresponding order for automatic cleaning.
Figure 7: Word list for automatic cleaning.
A.2 Manual Cleaning
Due to the specificity and complexity of language, manual cleaning remains an essential part of the process. To prevent virtual dialogue systems from exhibiting overly frequent anthropomorphic behavior, we identify instances of the Chinese word for "hug" (抱抱) and manually delete sentence snippets containing this term.
Appendix B Requirements for Dialogue Retention
B.1 Data Format
We provide the requirements for data format as follows:
  • 1.
    The generated conversations do not start with "求助者:" or "支持者:".
  • 2.
[…]
  • 3.
  • 4.
B.2 Dialogue Turns
Conversations comprising fewer than 5 turns will be discarded.
Appendix C Method
For brevity, we present the standard prompt and the standard prompt with a dialogue topic in Figure 8.
Figure 8: The standard prompt and the standard prompt with a dialogue topic. The two are separated by using the mark of "/" to differentiate between them. The key difference lies in whether a dialogue topic is introduced.
Appendix D Mechanism of Language Transformation
We propose two hypotheses: (1) When a single-turn dialogue is rewritten into a multi-turn dialogue, the similarity between the two is high (Attract). (2) When ChatGPT generates a multi-turn dialogue without introducing a single-turn dialogue, the similarity between the generated dialogue and a randomly selected single-turn dialogue is low (Repel). For better understanding, we present the mechanism of language transformation in Figure 9.
Figure 9: Mechanism for language transformation.
Appendix E Dialogue Topics Annotation
In this paper, to label the dialogue topics of generated dialogues, the hyperparameters of ChatGPT API during generation we used are set to the officially recommended default values, where temperature $\tau=1.0$ and $p=1.0$. Figure 10 shows the pipeline of dialogue topics annotation.
Figure 10: Pipeline of Dialogue Topics Annotation, where the content in bold is a placeholder.
Appendix F Dialogue Diversity
Metrics
@@ Appendix F Dialogue Diversity
F.1 Lexical Features
For lexical analysis, we utilize the ChatGLM2-6B tokenizer [^5] https://huggingface.co/THUDM/chatglm2-6b to tokenize the dialogue. To measure the lexical features, we adopt distinct-$n$ ($n=1,2,3$) metrics (Li et al., 2016), which are widely used for measuring the diversity of dialogue datasets. Each dialogue is first preprocessed into a single string without any speaker tokens. We provide statistics for 500 dialogues per prompt method, with PsyQA* serving as the reference point, detailed in Table 5.
Our proposed smile method results in rich vocabularies, with a significantly higher number of unique unigrams, bigrams and trigrams compared to the baseline methods. Specifically, a simple and fixed prompt tends to produce monotonous output. When dialogue topics are incorporated into a single, fixed prompt, the model’s output demonstrates substantial diversification. Furthermore, the smile method outperforms the baseline methods in terms of Distinct-1, Distinct-2 and Distinct-3.
F.2 Dialogue Topics
@@
To obtain dialogue topics for each dialogue in each prompt method, we design a prompt provided with 56 distinct dialogue topics, as illustrated in Appendix K and Figure 10, to automatically label dialogue topics for each dialogue with GPT-4. The model we use is gpt-4-0613. We present the information entropy for each prompt method in Table 6, demonstrating that the dialogues generated using the smile method are substantially more diverse than those generated using the standard method and are compatible with the stand.${}_{\mathrm{DT}}$ method, which uniformly samples dialogue topics.
Setting standard stand.${}_{\mathrm{DT}}$ smile
@@
Table 6: Information entropy of dialogue topics. The setting of limited topics indicates that the dialogue topics in the statistics are confined to the 56 dialogue topics we provide. The setting of unlimited topics denotes that the dialogue topics not only belong to our provided 56 distinct topics but also include topics generated spontaneously by GPT-4.
Appendix G Expert Evaluation
To illustrate the patterns of supporters’ strategies, we compile statistics on the flow patterns of these strategies, including "Supporting", "Challenging" and "Other".
G.1 Experimental Setup
PsyQA (Sun et al., 2021) contains 4012 QAs with supporters’ strategies in the original dataset. Xinling (Li et al., 2023) is a real-life multi-turn counseling dataset containing 550 dialogues with supporters’ strategies in the original dataset.
Three methods (standard, stand.${}_{\mathrm{DT}}$, and smile) each contain 500 dialogues, respectively. We use Xinling to train a classifier to predict the supporters’ strategies in dialogues generated with these three methods. The performance of the trained classifier is comparable to the results in the original paper.
G.2 Results
Supporters have three main strategies: "Supporting", "Challenging", and "Other", and we present the results in Table 7, 8 and 9. We highlight the gap greater the 0.2 in bold and the corresponding gap is in parentheses.
Strategy Supporting
Methods standard stand.DT SMILE PsyQA Xinling (Oracle)
Stage 1 0.854(+0.101) 0.847(+0.094) 0.615(-0.138) 0.474(-0.279) 0.753
Stage 2 0.547(-0.278) 0.520(-0.305) 0.432(-0.393) 0.312(-0.513) 0.825
Stage 3 0.419(-0.367) 0.365(-0.421) 0.360(-0.426) 0.179(-0.607) 0.786
Stage 4 0.468(-0.258) 0.387(-0.339) 0.390(-0.336) 0.107(-0.619) 0.726
Stage 5 0.808(+0.382) 0.774(+0.348) 0.662(+0.236) 0.154(-0.272) 0.426
Table 7: Pattern flow of "Supporting" strategy.
Strategy Challenging
Methods standard stand.DT SMILE PsyQA Xinling(Oracle)
Stage 1 0.140(+0.084) 0.150(+0.094) 0.372(+0.316) 0.348(+0.292) 0.056
Stage 2 0.452(+0.305) 0.478(+0.331) 0.565(+0.418) 0.598(+0.451) 0.147
Stage 3 0.576(+0.384) 0.632(+0.440) 0.636(+0.444) 0.752(+0.560) 0.192
Stage 4 0.516(+0.272) 0.603(+0.359) 0.586(+0.342) 0.847(+0.603) 0.244
Stage 5 0.112(-0.127) 0.157(-0.082) 0.231(-0.008) 0.794(+0.555) 0.239
Table 8: Pattern flow of "Challenging" strategy.
Strategy Other
Methods standard stand.DT SMILE PsyQA Xinling(Oracle)
Stage 1 0.006(-0.185) 0.003(-0.188) 0.013(-0.178) 0.178(-0.013) 0.191
Stage 2 0.001(-0.027) 0.002(-0.026) 0.003(-0.025) 0.090(+0.062) 0.028
Stage 3 0.005(-0.017) 0.003(-0.019) 0.004(-0.018) 0.069(+0.047) 0.022
Stage 4 0.016(-0.014) 0.010(-0.020) 0.024(-0.006) 0.046(+0.016) 0.03
Stage 5 0.080(-0.255) 0.069(-0.266) 0.107(-0.228) 0.052(-0.283) 0.335
Table 9: Pattern flow of "Other" strategy.
G.3 Conclusion
Dialogues generated by the three methods exhibit similar flow patterns of supporters’ strategies. Instructing ChatGPT to rewrite QAs into multi-turn dialogues will inevitably impact the strategy pattern in the generated dialogues. Dialogues generated by the three methods still exhibit significant gaps when compared to the Xingling (Oracle) dataset with respect to the flow patterns of supporters’ strategies.
Appendix H Implementation Details
Model Training.
We implement ChatGLM2-6B with the Transformers library. Table 10 shows the parameters of parameter-efficient fine-tuning.

### v3

@@ 4.2 Analysis
We propose two hypotheses: (1) When a single-turn dialogue is rewritten into a multi-turn dialogue, the similarity between the two is high (Attract). (2) When ChatGPT generates a multi-turn dialogue without introducing a single-turn dialogue, the similarity between the generated dialogue and a randomly selected single-turn dialogue is low (Repel). The mechanism of our proposed hypotheses is presented in Figure 3.
To assess the transformation feasibility of our SMILE method, we employ cosine similarity to calculate the transformation ratio. Specifically, a single-turn dialogue d is transformed into a multi-turn dialogue $\hat{d}$ using the SMILE method. A multi-turn dialogue generated by standard or standardT is denoted as $\bar{d}$. We can calculate cos($d$, $\hat{d}$) for the first hypothesis and cos($d$, $\bar{d}$) for the second hypothesis. We utilize the text embeddings obtained in Equation 3.
Figure 3: Mechanism for language transformation.
@@
Figure 4 presents the distribution of dialogue transformation among three methods. Our analysis concludes that single-turn dialogues can be successfully rewritten into multi-turn dialogues, ensuring that the dialogue generated by the proposed method is lifelike, rather than purely in the style of machine-generated text from ChatGPT.
Method
@@
5 Dialogue Diversity
To demonstrate the effectiveness of the SMILE method, we mainly focus on three aspects of diversity: lexical features, semantic features, and dialogue topics.
5.1 Lexical Features
For lexical analysis, we utilize the popular Chinese ChatGLM2-6B [^7] https://huggingface.co/THUDM/chatglm2-6b tokenizer that is widely used in NLP to tokenize the dialogue. To measure the lexical features, we adopt distinct-$n$ ($n=1,2,3$) metrics (Li et al., 2016), which are widely used for measuring the diversity of dialogue datasets. Each dialogue is first preprocessed into a single string without any speaker tokens. We provide statistics for 500 dialogues per prompt method, as presented in Table 1.
Our proposed SMILE method results in rich vocabularies, with a significantly higher number of unique unigrams, bigrams, and trigrams than the two baseline methods. Specifically, a simple and fixed prompt tends to produce monotonous output. The model output demonstrates substantial diversification when dialogue topics are incorporated into a single, fixed prompt. Furthermore, the SMILE method outperforms the baseline methods in terms of Distinct-1, Distinct-2, and Distinct-3.
5.2 Semantic Features
To measure the semantic diversity in a dialogue dataset, we propose to calculate the cosine similarity between every pair of different dialogues. This involves computing the pairwise cosine similarity for each pair of distinct dialogues, resulting in $\binom{500}{2}$ pairs and their corresponding cosine values, as described in Equation 3.
We present the results in Figure 5, which demonstrates that the median of the SMILE method is significantly lower than those of the baseline methods. The SMILE method exhibits the most extensive semantic diversity, aligning closely with the sampled dialogue of PsyQA. However, it is worth noting that the median of the SMILE method is more significant than that of PsyQA*. The reason behind this could be the introduction of token distribution from ChatGPT.
5.3 Dialogue Topics
@@
To obtain dialogue topics for each dialogue in each prompt method, we design a prompt provided with 60 distinct dialogue topics, as illustrated in Appendix G and Figure 11, to automatically label dialogue topics for each dialogue with Qwen1.5-110B-Chat (qwen). We present the information entropy for each prompt method in Table 2, demonstrating that the dialogues generated using the SMILE method are substantially more diverse than those generated using the standard method and are compatible with the standardT method, which uniformly samples dialogue topics.
Setting standard standardT SMILE
@@
Table 2: Information entropy of dialogue topics.
5.4 SMILECHAT Dataset
Through the comprehensive analysis of language transformation and dialogue diversity, we conclude that the proposed method can generate a lifelike and diverse chat dataset. Therefore, we utilize the SMILE method to guide ChatGPT in generating all multi-turn conversations based on PsyQA one round only, leading to a large-scale dialogue dataset.
Category Total Help-seeker Supporter
# Dialogues 55165 - -
# Utterances 1833856 693756 1140100
Turns per dialogue 5.7 - -
Utterances per dialogue 33.2 12.6 20.7
Avg. length per utterance 27.9 26.1 28.9
Table 3: Data statistics of the dialogue dataset, SMILECHAT.
Data Statistics
To ensure data quality, we impose stricter requirements on dialogue turns, retaining only dialogues with at least five turns. Thus, we compile a collection of 55165 conversations, SMILECHAT. Table 3 presents the statistics of the collected corpus.
Dialogue Exemplars
Multi-turn dialogue examples generated with the standard and standardT methods are illustrated in Figures 18 and 20. Further, a dialogue generated by the SMILE method is shown in Figure 22.
6 Dialogue System
We aim to build a high-quality multi-turn chat dataset for mental health support. It is non-trivial to evaluate the quality of a dialogue dataset, which is often assessed indirectly by the dialogue system. Therefore, we need to train a dialogue system and analyze its performance.
6.1 Mathematical Formulation
In order to train a dialogue system for mental health support, the first step is to split each full dialogue $d\sim\mathcal{D}$ into several training sessions. Specifically, a sampled $t$-turn dialogue session can be represented as $d_{t}=\left\{u_{1},r_{1},u_{2},r_{2},\dots,u_{t},r_{t}\right\}\sim\mathcal{D}$. Thus, we build a dialogue model that can predict the supporter’s utterance $r_{t}$ based on the dialogue history $h_{t}=\left\{u_{1},r_{1},u_{2},r_{2},\dots,u_{t}\right\}$. Our objective is to use our synthetic conversation dataset $\mathcal{D}$ to fine-tune a large language model $\pi_{0}$ using supervised learning, i.e., maximum likelihood estimates:
$J_{\mathrm{SFT}}(\theta)=\mathbb{E}_{(h_{t},r_{t})\sim\mathcal{D}}\left[\log\pi_{\theta}(r_{t}\mid h_{t})\right]$ (5)
where $\pi_{\theta}$ is initialized from $\pi_{0}$.
6.2 Experimental Setup
Backbone Model
To validate the dialogue quality of our collected dataset, we conduct a fine-tuning experiment on ChatGLM2-6B (Zeng et al., 2023).
Training Data
To meet the data format requirements for instruction-based fine-tuning, we split the dialogue into multiple sessions, with the supporter’s last utterance concluding each session. Additionally, we incorporate the system prompt (detailed in Appendix E) as a prefix to dialogue messages, following OpenAI’s data format.
Parameter-efficient Fine-tuning
To preserve the original capabilities of the model while adapting to downstream dialogue tasks and reducing computational costs, we employ Low-Rank Adaptation (LoRA, (Hu et al., 2021)) on all linear layers in the model for efficient fine-tuning.